A manufacturer’s serial numbers consist of 4 letters of the alphabet followed by 2 numerical digits. How many serial numbers are possible?
Answer by Hawksfan(61)

The number has the following format: LLLLNN
L = Letter
N = Number
There are 26 letters and 10 digits
So there would be 26*26*26*26*10*10 = 26^4 *100 = 456976